# Nick Stracke    Stefan Andreas Baumann11footnotemark: 1    Kolja Bauer11footnotemark: 1    Frank Fundel    Björn Ommer

### v1

@@ Acknowledgments
This project has been supported by the German Federal Ministry for Economic Affairs and Climate Action within the project “NXT GEN AI METHODS – Generative Methoden für Perzeption, Prädiktion und Planung”, the German Research Foundation (DFG) project 421703927, and the bidt project KLIMA-MEMES. The authors gratefully acknowledge the Gauss Center for Supercomputing for providing compute through the NIC on JUWELS at JSC and the HPC resources supplied by the Erlangen National High Performance Computing Center (NHR@FAU funded by DFG). Further, we would like to thank Owen Vincent for continuous technical support.
@@ C Additional Qualitative Samples
We provide additional qualitative samples for semantic correspondence matching in Figure 12 and for semantic segmentation in Figure 14.
@@
D Projection Heads
Projected Feature Maps
@@ D Projection Heads
Adapter Ablations
We show the remaining results of our ablation study from Sec. 4.6, including our investigation of the influence of different components of the projection head architecture and the influence of pre-training the projection heads. In our main configuration, we use a FiLM layer [38] in each FFN block to adaptively scale activations depending on the timestep $t$. We replace the FiLM layers with Adaptive RMS (AdaRMS) norm layers [57] and observe a performance degradation of $0.1$ percentage points for $\text{PCK}_{\text{bbox}}$. We conclude that removing the scale and shift information of the model’s feature by normalization is harmful and cannot be recovered by subsequent scaling and shifting.
E Evaluation Design Choices

### v2

@@ Acknowledgments
This project has been supported by the German Federal Ministry for Economic Affairs and Climate Action within the project “NXT GEN AI METHODS – Generative Methoden für Perzeption, Prädiktion und Planung”, the project “GeniusRobot” (01IS24083), funded by the Federal Ministry of Education and Research (BMBF), the bidt project KLIMA-MEMES, and the German Research Foundation (DFG) project 421703927. The authors gratefully acknowledge the Gauss Center for Supercomputing for providing compute through the NIC on JUWELS at JSC and the HPC resources supplied by the Erlangen National High Performance Computing Center (NHR@FAU funded by DFG project 440719683) under the NHR project JA-22883.
Further, we would like to thank Owen Vincent for continuous technical support.
@@ B Additional Quantitative Evaluations
Supervised Semantic Correspondence
We test our method in the supervised setting of TLFR [67] that uses an aggregation network [35, 66] to fuse feature maps. In this setting, our features achieve a $\text{PCK}_{\text{bbox}}\!$@$\alpha$=$0.1$ per image of $83.9$, outperforming SD2.1 features (83.2). This shows that even in a supervised setting, our features provide additional information that cannot be extracted by additional supervised training.
Distilled Text Conditioning
In our standard setting, we train CleanDIFT with image-text pairs since that is what the diffusion model expects as input. As a result, the model also requires a fitting text prompt for optimal feature performance at inference time. Here, we experiment with distilling the text conditioning during our fine-tuning to directly yield optimal features without the need for a prompt. To that end, we train a CleanDIFT version that depends neither on explicit nor implicit captions [66, 67] by distilling the text conditioning directly into its features. This results in a $1.2$ $\text{PCK}_{\text{bbox}}\!$@$\alpha$=$0.1$ gain over DIFT SD2.1 features with 8x ensembling and no text prompt when extracting features. Compared to simply training CleanDIFT without captions and no text conditioning during inference, the distillation improves performance by 0.4 $\text{PCK}_{\text{bbox}}\!$@$\alpha$=$0.1$.
@@ C Additional Qualitative Samples
We provide additional qualitative samples for semantic correspondence matching in Figure 12 and for semantic segmentation in Figure 13. Figure 16 shows a PCA visualization of our features revealing that they are less noisy than standard features.
D Depth Probes Analysis
@@
SPair71k (Test) COYO (Generic, Ours) ImageNet (mismatched) 61.42 61.43 60.78
Table 6: SPair $\text{PCK}_{\text{bbox}}\!$@$\alpha$=$0.1$ when training on different datasets. A generic dataset performs best and training on the test dataset does not yield any additional gains.
E Additional Projection Head Ablations
We investigate the influence of different components of the projection head architecture and the influence of pre-training the projection heads following the setup in  Sec. 4.6. In our main configuration, we use a FiLM layer [45] in each FFN block to adaptively scale activations depending on the timestep $t$. We replace the FiLM layers with Adaptive RMS (AdaRMS) norm layers [65] and observe a performance degradation of $0.1$ percentage points for $\text{PCK}_{\text{bbox}}$. We conclude that removing the scale and shift information of the model’s feature by normalization is harmful and cannot be recovered by subsequent scaling and shifting.
@@ E Additional Projection Head Ablations
F Other Diffusion Backbones
We evaluate whether our method applies to other diffusion backbones by evaluating semantic correspondence performance in a DIFT [62] setting and find gains in all cases (cf. Tab. 7). This shows that diffusion models suffer from noisy features independent of their size and architecture, and that CleanDIFT can successfully remedy that. Note that for SDXL (Turbo) and Flux we trained LoRAs [26] instead of fully fine-tuning the entire model due to their large model size. We adapt the self-attention and FFN layers with LoRAs of rank 64. We swept all models for the optimal timestep and feature map and used that for the PCK calculation.
Backbone PCK@$\alpha$ Gain ()($\uparrow$)( ↑ ) $\alpha_{\mathrm{img}}$ = 0.1 $\alpha_{\mathrm{bbox}}$ = 0.1 SDXL [46] 1.7 1.6 SDXL Turbo [55] 3.2 3.7 PIXART-$\alpha$ [7] 2.7 2.2 Flux [14] 9.4 8.1
Table 7: We assess the effectiveness of our features for other diffusion backbones such as a much larger UNet (SDXL), diffusion transformers (PIXART-$\alpha$, Flux), and a distilled model (SDXL Turbo). The PCK quantifies the gain obtained by using our features for semantic correspondences using the standard DIFT [62] setup when compared to the standard features. We show that using our CleanDIFT features leads to better performance for all backbones.
G Dataset Considerations
We evaluated different datasets and find simple choices to suffice. Specifically, a random (only filtered to a minimum size of $768^{2}$) subset of COYO-700M with $\sim$3k images sufficed for our optimal results. COYO-700M is a dataset that is close to the original model’s pretraining setting and not tailored to any of the categories relevant to our downstream evals. We show zero-shot semantic correspondence results across different fine-tuning datasets in Tab. 6. Overfitting on the test setting, e.g., by training on SPair71k while also evaluating on that dataset does not give gains over our non-tailored version, but using a dataset that has little overlap with the distribution of the target task (e.g., ImageNet) results in reduced performance.
Image DIFT [62] Ours Refer to caption
Figure 16: Principal component visualization of our features. One 4-component feature PCA is computed per column, with the first component being thresholded at 0 to obtain the foreground and the remaining being mapped to RGB color. CleanDIFT produces similarly semantically useful features as DIFT, while exhibiting less noise.
H Evaluation Design Choices